OpenStudy (anonymous):

What is the graph of the system? y ≤ -x – 1 y ≥ 2x + 4

OpenStudy (anonymous):

This is quite a big assignment eh?

OpenStudy (anonymous):

ive already tried doing this somehow i got this wrong

OpenStudy (anonymous):

am I correct reading it as y less than or equal to -x -1 AND y greater than or equal to 2x+4

OpenStudy (anonymous):

these are the questions i got wrong i missed 9 out of 20

OpenStudy (anonymous):

> less than <greater than

OpenStudy (anonymous):

always read thes symbols from left to right . < means that the left is smaller than the right

OpenStudy (anonymous):

now im confused

OpenStudy (anonymous):

For example 3 < 4 is read as 3 is less than 4.

OpenStudy (anonymous):

so im wrong, ehhhhh im lost

OpenStudy (anonymous):

which bit is scaring you most?

OpenStudy (anonymous):

the part where ive had my signs backwards now im lost

OpenStudy (anonymous):

so i have the points graphed just dont know what to shade in lol

OpenStudy (anonymous):

This may help you or it may give you an idea. I googled "Graphing inequalities" http://www.purplemath.com/modules/ineqgrph.htm Dont give up

OpenStudy (anonymous):

I always graph the line and the after i look at my graph i choos a point. BTW do you colour in the region that makes the inequality true?

OpenStudy (anonymous):

hahahahah your probly going to say wow this chick is stupid but those pages didnt help me at all

OpenStudy (anonymous):

honestly my brain is still stuck on 3rd grade math lol

OpenStudy (anonymous):

No I did not know you were a chick !

OpenStudy (anonymous):

Some pages don't give instant answers and Khan academy suff is excellent if you give it time.

OpenStudy (anonymous):

wow the time limit is almost up and i have 6 more questions to answer if i dont answer them by the end of the time limit then it automatically subbmits its self and i fail again

OpenStudy (anonymous):

i guess ill have to guess on this question again

OpenStudy (anonymous):

Can you redo the test at a later date? You will know what you need to learn for next time.

OpenStudy (anonymous):

no i only get to tries

OpenStudy (anonymous):

2

OpenStudy (anonymous):

That makes it tough!

OpenStudy (anonymous):

yep and ive already failed it once

OpenStudy (anonymous):

  The first one is for y<= -x-1 and the second is for y>=2x+4 How to sketch: 1. Sketch the the line (eg. y<= -x-1 becomes y=-x-1) 2. if it is less than (<) then shade in on the left side. If it is greater than (>) shade in on the right side 3. if it greater than/less than, then you don't darken the line, if it is equal to and less than equal to, then darken the line